Steve powered up and cooled the wavefront sensor in the afternoon. no issues

0:50 OSCO HD19332
  • 0.211 optical gain applied TN#10839, TN#11137
  • NCPA on, TN#011712 Safe skipping, gains turned down a touch, seeing 1.7" on the guider (corrected). DIMM 2.3" but steve is still focusing it. TN#012058
6:44 Preset to Science field ESM left side, SL right side
  • Magntiude set to 14.51
  • Estimated magnitude on WFS 15.7, 6.6 photons/subapp, 39 counts average pupil intensity
  • bin4 400 Hz 0.59 0.34 0.10
  • 1.1" on luci2 guider TN#015647
  • Failed to reclose after 50" offset. Manually reclosed, safeskipped for a few frames with LNGains until recentered then applied gains.
  • 0.59" on LUCI1 and 0.96" on LUCI2 images
10:00 M3 selector issue
  • The m3 was not recoverable. Moving to MODS. WFS has been warmed and powered off.

-- JenniferPower - 02 Jan 2021
Topic revision: r2 - 02 Jan 2021, JenniferPower
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback